Two different forecasting techniques (F1 and F2) were used to forecast demand for cases of bottled water. Actual demand and the two sets of forecasts are as follows:

PREDICTED DEMAND
Period Demand F1 F2
1 68 62 67
2 75 69 68
3 70 72 70
4 74 66 70
5 69 73 72
6 72 65 76
7 80 74 79
8 78 74 85


a. Compute MAD for each set of forecasts. Given your results, which forecast appears to be more accurate? (Round your answers to 2 decimal place.)

MAD F1
MAD F2


  (Click to select)   F1   F2   None  appears to be more accurate.

b. Compute the MSE for each set of forecasts. Given your results, which forecast appears to be more accurate? (Round your answers to 2 decimal places.)

MSE F1
MSE F2


  (Click to select)   F1   F2   None  appears to be more accurate.

c. In practice, either MAD or MSE would be employed to compute forecast errors. What factors might lead a manager to choose one rather than the other?

Either one might already be in use, familiar to users, and have past values for comparison. If     (Click to select)   control charts   tracking signals  are used, MSE would be natural; if   (Click to select)   tracking signals   control charts  are used, MAD would be more natural.

d. Compute MAPE for each data set. Which forecast appears to be more accurate? (Round your intermediate calculations to 2 decimal places and and final answers to 2 decimal places.)

MAPE F1
MAPE F2

  (Click to select)   F1   F2   None  appears to be more accurate.
Answer #1

FORECAST 1


PERIOD

ACTUAL

FORECAST

DEVIATIONS

ABSOLUTE DEVIATION

ABS DEV / DEMAND * 100

DEVIATION^2

1

68

62

6

6

8.82

36

2

75

69

6

6

8

36

3

70

72

-2

2

2.86

4

4

74

66

8

8

10.81

64

5

69

73

-4

4

5.8

16

6

72

65

7

7

9.72

49

7

80

74

6

6

7.5

36

8

78

74

4

4

5.13

16

SIGMA

31

43

58.64

257


MAD = SIGMA(ABS DEV) / N

MAD = 43 / 8 = 5.38

MSE = SIGMA(DEV^2) / N

MSE = 257 / 8 = 32.13

MAPE = SIGMA((ABS DEV / ACTUAL DEMAND * 100) / N

MAPE = 58.64 / 8 = 7.33


FORECAST 2


PERIOD

ACTUAL

FORECAST

DEVIATIONS

ABSOLUTE DEVIATION

ABS DEV / DEMAND * 100

DEVIATION^2

1

68

67

1

1

1.47

1

2

75

68

7

7

9.33

49

3

70

70

0

0

0

0

4

74

70

4

4

5.41

16

5

69

72

-3

3

4.35

9

6

72

76

-4

4

5.56

16

7

80

79

1

1

1.25

1

8

78

85

-7

7

8.97

49

SIGMA

-1

27

36.34

141

MAD = SIGMA(ABS DEV) / N

MAD = 27 / 8 = 3.38

MSE = SIGMA(DEV^2) / N

MSE = 141 / 8 = 17.63

MAPE = SIGMA((ABS DEV / ACTUAL DEMAND * 100) / N

MAPE = 36.34 / 8 = 4.54



2. FORECAST 2 APPEARS TO BE BETTER IN ALL THE THREE CASES